OpenStudy (anonymous):

Given that the intergral from 0 to 1 of 3x*sqrt(x^2+4)dx=5*sqrt(5)-8, what is the intergral from 0 to 1 of 3u*sqrt(u^2+4)du?

OpenStudy (anonymous):

in the definite integral \[\int\limits_{a}^{b}f \left( x \right)dx=\int\limits_{a}^{b}f \left( y \right)dy\]

OpenStudy (anonymous):

So the answer would be \[5\sqrt{5}-8\] as well?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

yw